# Enhancing Large Language Models for Text-to-Testcase Generation

### v1

@@ Abstract
Context: Test-driven development (TDD) is a widely employed software development practice that involves developing test cases based on requirements prior to writing the code. Although various methods for automated test case generation have been proposed, they are not specifically tailored for TDD, where requirements instead of code serve as input.
Objective: In this paper, we introduce a text-to-testcase generation approach based on a large language model (GPT-3.5) that is fine-tuned on our curated dataset with an effective prompt design.
Method: Our approach involves enhancing the capabilities of basic GPT-3.5 for text-to-testcase generation task that is fine-tuned on our curated dataset with an effective prompting design. We evaluated the effectiveness of our approach using a span of five large-scale open-source software projects.
Results: Our approach generated 7k test cases for open source projects, achieving 78.5% syntactic correctness, 67.09% requirement alignment, and 61.7% code coverage, which substantially outperforms all other LLMs (basic GPT-3.5, Bloom, and CodeT5). In addition, our ablation study demonstrates the substantial performance improvement of the fine-tuning and prompting components of the GPT-3.5 model.
@@ 1 Introduction
Test-Driven Development (TDD) is an agile software development methodology that prioritizes creating test cases before the actual coding process (Beck, 2022). This methodology is structured around iterative cycles that involve the development of test cases based on a given requirement, the implementation of code to pass these test cases, and then refactoring to continuously improve the code. TDD is claimed to significantly improve code quality and handle requirements better than ’more traditional, heavy weight predictive methods’ (Wasmus and Gross, 2007). The implementation of TDD led to a reduction in defects by approximately 50%, in comparison to a similar system developed with ad-hoc unit testing methods (Maximilien and Williams, 2003). Despite these benefits, one of the primary drawbacks of this methodology is the time and effort required to write test cases. The initial investment in creating comprehensive and effective test cases is substantial and represents the largest overhead cost in the TDD process (Beck, 2022). However, this investment is a critical component of TDD.
@@
The Large Language Models (LLMs), being advanced language models, are capable of performing various Natural Language Processing (NLP) tasks such as text generation, question answering, and classification. Specifically in the field of software engineering (SE), LLMs are increasingly being applied for tasks such as requirements engineering, code generation, vulnerability identification, test generation, and software maintenance. The use of LLMs for generating test cases from natural language descriptions of software methods (text-to-test case) is a relatively new application.
@@
In this paper, we introduce a Text-to-Test generation approach, which involves a fine tuning component for the basic GPT-3.5 using our dataset specifically curated from open-source Java projects and an effective prompting design. The model fine-tuning is aimed to teach the LLM the mapping between the method description (text), and the corresponding test (testcase). Once the basic GPT-3.5 model is fine-tuned, during the model inference phase, an unseen method description (i.e., the method description is referred to as the requirement) with the designed prompt is used as input into the fine-tuned GPT-3.5 to generate the corresponding test case without requiring the code. Crucially, our approach differs from traditional methods by generating test cases directly from method descriptions without requiring the implemented code (method). This allows for the generation of test cases in the absence of actual code, a significant shift from existing practices.
@@
We compare our approach with basic GPT-3.5 (OpenAI, 2023), Bloom (Luccioni et al., 2022), and CodeT5 (Wang et al., 2021) fine-tuned on our newly curated benchmark dataset, consisting of 163k pairs of method descriptions and test cases. We developed an evaluation dataset with 7K test methods that span five large-scale open source software projects Apache Commons Lang (Lang, 2022), JFreeChart (JFreeChart, 2022), Apache Common CLI (CLI, 2022), Apache Common CSV (CSV, 2022), Google Gson (Gson, 2019). For each model, we evaluate the quality of the generated test cases in three aspects: syntax correctness, requirement alignments, and code coverage. In summary, our objective is to address the following research questions (RQs):
RQ1. How effective is the proposed approach for the Text-to-Test generation task?
@@ Results
.  The proposed approach generates test cases that are 78.5% syntactically correct and 67.09% aligned with the requirements while achieving a code coverage of 61.7%, which substantially outperforms the baselines (basic GPT-3.5-turbo, Bloom, and CodeT5).
RQ2. What is the effect of the fine-tuning step for the LLM-based text-to-testcase generation task?
@@
.  The proposed fine-tuning component improves syntax correctness by 223%, requirements alignment by 164%, and code coverage by 153%.
RQ3. What is the effect of the prompting step for the LLM-based text-to-testcase generation task?
Results
.  The proposed prompt design improves the syntax correctness by 124%, the requirements alignment by 82%, and the code coverage by 58%.
RQ4. What types of errors occur in the incorrectly generated test cases?
@@
Based on our results, we draw the following recommendations:
  • 1.
    Recommendation 1: The basic usage of ChatGPT may not produce a desirable performance for the text-to-testcase generation task. Our RQ1 shows that the basic usage of ChatGPT (GPT-3.5-turbo with no fine-tuning and a basic prompt) may not achieve a desirable performance as we observe that it performs worse than other LLMs like Bloom.
  • 2.
    Recommendation 2: Both fine-tuning and effective prompting design should always be considered when using a large language model. Our RQ1 shows that our proposed approach performs the best for the text-to-testcase generation task, while our RQ2 and RQ3 confirm the substantial benefits of the fine-tuning and the prompting component. However, both components should be considered together, as the results show that both fine-tuning and effective prompting designs outperform LLMs with one individual component alone.
@@
  1. 1.
    We present a new Text-to-Test generation approach based on a fine-tuned large language model with an effective prompting design.
  2. 2.
    The ablation study confirms the substantial benefits of fine-tuning and prompt design for the text-to-test generation task.
@@
Structure. Section 2 discusses background and related work. Section 3 presents our approach. Section 4 presents our experiment. Section 5 presents the results. Section 6 discloses the threats to validity, and Section 7 concludes the paper.
@@ 2.1 Test Driven Development (TDD)
Test Driven Development (TDD) (Beck, 2022; Astels, 2003) is a software development practice that suggests writing test cases before implementing source code. Starting with a comprehensive requirement, TDD follows an iterative process of providing a small unit test case that aligns with the requirement while developing a small chunk of source code to pass the test case at a time. The practice aligns with the Agile Development Methodology that emphasizes fast continuous feedback loops and enhances collaboration between team members. Recent research found that TDD can increase the understanding of requirements and produce a higher number of test cases than the traditional code-first method (Mueller and Zeller, 2002; Janzen, 2008; Erdogmus and Arisan, 2005). Despite the numerous benefits discussed in the literature, TDD is still manual, time-consuming, and labor-intensive, highlighting the need for automated Text-To-Testcase generation.
@@ 2.2 Large Language Models (LLMs)
A language model (LM) is a statistical model trained to predict the next word in a sequence (Jurafsky and Martin, 2020), e.g., for the sequence “Paris is the capital of”, it will predict with a very high likelihood that “France” is the next word in the sequence. Large language models (LLMs) are LMs with a substantially large number of weights and parameters and a complex training architecture trained to perform various downstream NLP tasks, e.g., text generation, question answering, and text classification (Wei et al., 2022). The recent advancements in LLMs, most notably with chatbot platforms built on LLMs, e.g., OpenAI’s ChatGPT, have led to several applications in software engineering (SE) (Nguyen-Duc et al., 2023; Hou et al., 2023). These include, among others, requirements engineering automation tasks (Arora et al., 2023), code generation and completion (Fan et al., 2023), software vulnerability (Fu et al., 2023), test generation (Yuan et al., 2023), and software maintenance (Nguyen-Duc et al., 2023).
@@
Prompting. In generative AI models (built on LLMs), the input is provided as a natural language instruction to elicit a desired response or output (Brown et al., 2020). The input is known as a “prompt”. How well a prompt is crafted determines the quality and relevance of the AI’s output. There are different prompting strategies, and one typically needs to experiment with various prompts to elicit accurate responses or outputs from the models (Nguyen-Duc et al., 2023).
Fine-tuning In the context of Generative AI and ML refers to taking a pre-trained model and further training it on a specific, usually smaller, dataset to adapt it to a particular task or set of tasks. This process is prevalent in applications where a model, initially trained on a diverse and general dataset, needs to be specialized for more specific purposes. This specialized dataset is tailored to meet the particular needs or context of the intended application or task. This approach is beneficial as it leverages the extensive learning from the initial training phase, making the fine-tuning process more efficient than training a new model from scratch on the specialized data. Therefore, fine-tuning enhances the model’s performance and accuracy in tasks requiring a more focused understanding or specialized knowledge (Brown et al., 2020).
@@ 2.3.1 Automated Test Case Generation
Automated test case generation is an emerging technique that leverages neural networks and other deep learning architectures to automatically create test cases for software applications. This approach aims to enhance the efficiency and effectiveness of software testing by automating the generation of test inputs and expected outputs. Researchers leveraged various approaches for automated test case generation, e.g., randomization, heuristics, search-based algorithms, and deep learning.
Random-based test case generation approaches like Randoop Pacheco and Ernst (2007) are proposed to automatically generate test cases randomly. Randoop uses a feedback-directed random test generation algorithm. It starts by generating a sequence of method calls and this sequence is generated randomly. Randoop uses feedback from previous tests to guide the selection process. For example, Randoop may be more likely to generate a sequence of method calls that includes methods that have not been tested in the past. It observes the results of the execution to determine whether the program behaved as expected and creates test cases based on the results of the execution.
Search-based test case generation approaches like Evosuite Fraser and Arcuri (2011) uses an evolutionary algorithm to generate test cases that maximize code coverage. It works by first generating a random population of test cases. Then, it executes the test cases and measures the code coverage. The test cases that achieve the highest code coverage are selected to be the parents of the next generation of test cases. EvoSuite mutates the parents to create the next generation of test cases. The mutations are designed to increase the code coverage of the test cases. It repeats this process until it reaches a time limit or until it finds a set of test cases that achieve a certain level of code coverage.
@@ 2.3.2 LLMs for Test Case Generation
Large Language Models (LLMs) have emerged as powerful tools in various fields of natural language understanding and generation, including software engineering (Hou et al., 2023). Notably, Hou et al. (Hou et al., 2023) and Wang et al. (Wang et al., 2023) found that LLMs have been widely used to support various software testing tasks. For example, Schafer et al. (Schäfer et al., 2023) proposed TestPilot, a LLM-based test generation tool for JavaScript that automatically generates unit tests for all API functions in an npm package. Siddiq et al. (Siddiq et al., 2023) leveraged Large Language Models (LLMs) for test case generation. Similarly, Yuan et al. (Yuan et al., 2023) and Xie et al. (Xie et al., 2023) leveraged ChatGPT for test case generation.
@@
3 Enhancing Large Language Models for Text-To-Test Generation
Problem Formulation. In this paper, we propose a text-to-testcase generation approach based on a fine-tuned LLM with an effective prompting design. In particular, we formulate the text-to-testcase as a generation task using a fine-tuned LLM (GPT-3.5-turbo). The model is fine-tuned on a large dataset, and is mined and curated from open-source Java projects. During the model fine-tuning, our LLM aims to learn the mapping between the method description (text), which is the comments or documentation that describe the purpose, functionality, and usage of a method, and the corresponding test case (testcase) which is a test method that includes a set of assert statements designed to verify whether the implemented method (method) is aligned with the given requirement or not (i.e., functionally correct). Once fine-tuned, during the model inference phase, an unseen method description is used as input into the fine-tuned LLM to generate the corresponding test cases without requiring the implemented method (method), which differs from most existing test case generation approaches. Our approach consists of three phases: (1) model fine-tuning, (2) prompt design, and (3) model inference. Below, we describe the details of our approach (Figure 1).
3.1 Model Fine-tuning
While LLMs demonstrate impressive capabilities in understanding and generating natural language (Hou et al., 2023), they may lack the nuanced contextual understanding required for specialized applications, e.g., the text-to-testcase generation task. The fine-tuning of LLMs is essential to enhance their performance and adaptability to our text-to-testcase generation task. Thus, it is important to have a representative and well-labeled dataset for the specific domain to achieve effective fine-tuning results. However, the existing test case generation datasets are not suitable for our task. For example, Tufano et al. (Tufano et al., 2022) introduced the Method2Test dataset, which is a pair of <<<Method,Testcase>>>, but lacking the corresponding method description that is required for our approach. Thus, we need to carefully curate the fine-tuning dataset.
3.1.1 Dataset Curation
For the fine-tuning of the model, we started from the Methods2Test dataset (Tufano et al., 2022). This dataset consists of 780K pairs of Java focal methods and their corresponding test cases, collected from 91K open-source Java projects. Since the method description is unavailable, we extended the Methods2Test dataset. Starting from the 91K open-source Java projects, we cloned each project from the provided GitHub URL and performed the following steps:
(Step 1) Collecting Code Structural Information. The structural information of the code, such as method and class names are very important as we used to identify the relationship between the focal methods (i.e., methods to be tested) and their corresponding test methods. To collect such information, we used the tree-sitter parser (Tree-sitter, 2022) to analyze the structure of source code within the given project based on the detailed abstract syntax tree (AST) representation of source code. With this step, we are able to identify the list of method and class names within the project.
(Step 4) Identify Test Methods and Focal Methods. We identify the test methods for each test class based on the @Test annotation. Similar to Step 3, we apply the method name matching heuristic (i.e., removing Test prefix or suffix in the method name) to identify the corresponding focal method and its method description.
With these steps, we construct the dataset <<<Text,Testcase,Method>>> in the format of triplets. The dataset consists of 163k pairs of JUnit test cases and corresponding method description. We note that the methods will not be used for model fine-tuning like prior work, but are only used for evaluating test cases.
3.1.2 Fine-Tuning
When performing model fine-tuning, each training example generally consists of a single input (i.e., method description) example and its associated output (i.e., test case) without giving detailed instructions or including multiple examples in the same prompt normally used for the base models (GPT-3.5). Following the OpenAI’s legacy fine-tuning best practices [^1] https://platform.openai.com/docs/guides/legacy-fine-tuning, each training example typically consists of a single input example and its associated output, which is typically constructed in the following format: {"input": "<Text>", "output": "<test case>"}. We extract method descriptions and corresponding test cases into a list of prompts and completions for fine-tuning. We initiate a fine-tuning job using OpenAI’s CreateFineTuneJob.create method, specifying the training data and model to fine-tune (GPT-3.5-turbo). We use ChatCompletion.create to generate test cases with the fine-tuned model, providing: The model engine ID from the fine-tuning job, the input is the method description prompt. The output is completion of the generated test case.The template includes the {"prompt": "<Text>", "completion": "<test case>". The input is a single input-output pair. We used the curated dataset constructed in Section 3.1.1 for model fine-tuning (the pairs of <text, test case>). Once completed, we use the fine-tuned model for generating test cases in the next step.
Figure 2: Structure of our Improved prompt and Basic prompt.
3.2 Model Inference
3.2.1 Test Case Generation
For test case generation, the most important step was to develop a prompt that works well, as the prompts provide context and guidance to the model. Without a (clear) prompt, the model would not know what information or response is expected. The process of formulating prompts for software engineering tasks is inherently subjective, and there are no established empirical guidelines for optimizing prompt effectiveness and robustness. We conducted several experiments involving various prompt combinations, ranging from basic to customized prompts, to address concerns associated with prompt optimization and devised our improved prompt. We present two prompts below. A basic prompt (our starting point) and an improved prompt (after several iterations).
Basic prompt It is a straightforward instruction for LLM that does not include additional instructions, context, or specialized formatting. A generic or minimal input relies on the model’s capabilities to generate a response based on the provided context. We use the basic prompt as a context in the model. The structure of the basic prompt is as follows:
"/* Write a junit test case /**Description **/"
Improved prompt It is the iterated version based on multi-step optimizations. We tested the prompt in the first step on a few samples from a small validation set (not from our curated or evaluation set to avoid data leakage), manually analyzed the generation performance, and optimized the generation results by providing ChatGPT with a series of new prompts.
As in Figure 2, we designed the prompt to instruct the LLM with a more specific context. It emphasizes the importance of proper and relevant assert statements, avoiding repetition, following a structured test case format, and adding the context details for a more structured test case. A detailed breakdown of the prompt template components and their intended purpose is:
  • 1.
    Introductory Comment: Instructs the model to focus on creating a test case, highlighting the need for relevant assert statements and avoiding repetition.
  • 2.
  • 3.
    Class and Method Name: Specifies the target class and method name for which the test is to be written, ensuring the generated test is relevant.
  • 4.
    Method Description: Provides a description of the method, giving context to generate a test case.
3.2.2 Post-Processing
When extracting tests from LLM’s response, syntactic errors may occur because the tokens limit is exceeded, resulting in truncated tests. We thus perform a three-part post-processing on the generated test cases.
  1. 1.
  2. 2.
  3. 3.
    Correcting Incomplete Parentheses: We noticed that some of the generated test cases contain incomplete parentheses, such as () or . To address this issue, we perform corrections on these long sequence test cases, ensuring that all parentheses are correctly balanced.

### v2

@@ Abstract
Context: Test-driven development (TDD) is a widely employed software development practice that involves developing test cases based on requirements prior to writing the code. Traditional TDD is time-consuming, since writing test cases based on a given requirement is a complex task. Although various methods for automated test case generation have been proposed, they are not specifically tailored for TDD, where requirements instead of code serve as input.
Objective: In this paper, we introduce a text-to-testcase generation approach based on a large language model (GPT-3.5) that is fine-tuned on our curated dataset with an effective prompt design. We aim to investigate the performance of LLMs-based text-to-testcase generation task based on two contexts, i.e., when LLMs are leveraged by fine-tuning and prompting.
Method: Our approach involves enhancing the capabilities of basic GPT-3.5 for text-to-testcase generation task that is fine-tuned on our curated dataset with an effective prompting design. We evaluated the effectiveness of our approach using a span of five large-scale open-source software projects. In total, we investigate 32 variations of eight LLMs (i.e., basic GPT-3.5-turbo, Incoder, Starcoder, CodeT5, Bloom, CodeGemma, CodeLlama, and Gemini).
Results: Our approach generated 7k test cases for open source projects, achieving 78.5% syntactic correctness, 67.09% requirement alignment, 61.7% code coverage, and 18.9% mutation score which substantially outperforms all other LLMs. Furthermore, our ablation study demonstrates the substantial performance improvement of the fine-tuning and prompting components of the GPT-3.5 model.
@@ 1 Introduction
Test-Driven Development (TDD) is a commonly used software development practice [1, 2, 3] for designing and developing software, widely adopted by agile software development teams that prioritize creating test cases before the actual coding process. This methodology is structured around iterative cycles that involve the development of test cases based on a given requirement, the implementation of code to pass these test cases, and then refactoring to continuously improve the code. TDD is claimed to significantly improve code quality and handle requirements better than ’more traditional, heavy weight predictive methods’ [4]. The implementation of TDD led to a reduction in defects by approximately 50%, in comparison to a similar system developed with ad-hoc unit testing methods [5]. TDD empowers programmers to enhance code quality both before and after writing code. By regularly assessing testability and refactoring it to improve without executing the program, developers can ensure that production code is always test-ready. Despite these benefits, one of the primary drawbacks of this methodology is the time and effort required to write test cases. The initial investment in creating comprehensive and effective test cases is substantial and represents the largest overhead cost in the TDD process [6]. However, this investment is a critical component of TDD.
@@
The Large Language Models (LLMs), being advanced language models, are capable of performing various Natural Language Processing (NLP) tasks such as text generation, question answering, and classification. Specifically in the field of software engineering (SE), LLMs are increasingly being applied for tasks such as requirements engineering, code generation, vulnerability identification, test generation, and software maintenance. Large Language Models (LLMs) have significantly excelled in automatic code generation [11, 12], code completion [13] , automatic patch generation  [14]  [15], comment generation  [16], and many others  [17]. While LLMs demonstrate proficiency in various automated programming tasks, their application in automatic test case generation, particularly under Test-Driven Development (TDD) conditions, remains limited. To the best of our knowledge, there is a lack of in-depth investigation into the performance of LLMs for Text-to-Testcase generation tasks.
While other automatic test case generation approaches are proposed, none of these approaches are designed for TDD (i.e., taking requirement as input), are not able to generate test cases without code (i.e., mostly they take code as input). Therefore, these limitations highlight the need for an automated test case generation approach that takes a requirement as input and can generate alternative test cases beyond the ones that exactly match the ground truth test cases.
As a result, it remains an open question whether LLMs can be effectively utilized for test case generation from textual descriptions. To the best of our knowledge, there is a lack of in-depth investigation into the performance of LLMs for Text-to-Testcase generation tasks. In this work, we aim to fill this gap by exploring how LLMs perform on Text-to-Testcase generation tasks.
@@
In this paper, we introduce a Text-to-Testcase generation approach, which involves a fine tuning component for the basic GPT-3.5 using our dataset specifically curated from open-source Java projects and an effective prompting design. The model fine-tuning is aimed to teach the LLM the mapping between the method description (text), and the corresponding test (testcase). Once the basic GPT-3.5 model is fine-tuned, during the model inference phase, an unseen method description (i.e., the method description is referred to as the requirement) with the designed prompt is used as input into the fine-tuned GPT-3.5 to generate the corresponding test case without requiring the code. Crucially, our approach differs from traditional methods by generating test cases directly from method descriptions without requiring the implemented code (method). This allows for the generation of test cases in the absence of actual code, a significant shift from existing practices.
@@
In this work, we aim to investigate the performance of LLMs-based text-to-testcase generation task based on two contexts, i.e., when LLMs are leveraged by fine-tuning and prompting. In particular, we evaluate eight LLMs (i.e., GPT-3.5 [18], Bloom [19], CodeT5 [20], Incoder [21], Starcoder [22], CodeGemma [23], CodeLlama [12], and Gemini [24]) fine-tuned on our newly curated benchmark dataset, consisting of 163k pairs of method descriptions and test cases. We developed an evaluation dataset with 7K test methods that span five large-scale open source software projects Apache Commons Lang [25], JFreeChart [26], Apache Common CLI [27], Apache Common CSV [28], Google Gson [29]. For each model, we evaluate the quality of the generated test cases in four aspects: syntax correctness, requirement alignments, code coverage and mutation score. In summary, our objective is to address the following research questions (RQs):
RQ1. How effective is the proposed approach for the Text-to-Testcase generation task?
@@ Results
.  The proposed approach generates test cases that are 78.5% syntactically correct and 67.09% aligned with the requirements while achieving a code coverage of 61.7%, and 18.9% mutation score which substantially outperforms the baselines (basic GPT-3.5-turbo, Incoder, Starcoder, CodeT5, Bloom, CodeGemma, CodeLlama, and Gemini).
RQ2. What is the effect of the fine-tuning and prompting step for the LLM-based text-to-testcase generation task?
@@
.  The proposed fine-tuning component improves syntax correctness by 223%, requirements alignment by 164%, the coverage of the code by 153%, and mutation score improved by 273%. The proposed prompt design improves the syntax correctness by 124%, the requirements alignment by 82%, code coverage by 58%, and mutation score improved by 150%.
RQ3. What types of errors occur in the incorrectly generated test cases?
@@
Based on our results, we draw the following recommendation:
The basic usage of ChatGPT may not produce desirable performance for the text-to-testcase generation task. Both fine-tuning and effective prompting design should always be considered when using a large language model. Our RQ1 shows that our proposed approach performs the best for the text-to-testcase generation task, while our RQ2 confirms the substantial benefits of the fine-tuning and the prompting component. The results show that both fine-tuning and effective prompting designs outperform LLMs with one single individual component alone.
@@
  1. a)
    We introduce a novel approach that utilizes Large Language Models (LLMs) for Text-to-Testcase generation based on fine-tuning with an effective prompting design. We are the first to investigate the performance of LLMs-based for the Text-to-Testcase generation task.
  2. b)
    The ablation study confirms the substantial benefits of fine-tuning and prompt design for the Text-to-Testcase generation task.
@@
Structure. Section 2 discusses background and related work. Section 3 presents our experiment design. Section 4 presents our approach. Section 5 presents the results. Section 6 presents the discussion. Section 7 discloses the threats to validity, and Section 8 concludes the paper.
@@ 2.1 Test Driven Development (TDD)
Test Driven Development (TDD) [6, 30] is a software development practice that suggests writing test cases before implementing source code. Starting with a comprehensive requirement, TDD follows an iterative process of providing a small unit test case that aligns with the requirement while developing a small chunk of source code to pass the test case at a time. The practice aligns with the Agile Development Methodology that emphasizes fast continuous feedback loops and enhances collaboration between team members. Recent research found that TDD can increase the understanding of requirements and produce a higher number of test cases than the traditional code-first method [31, 32, 33]. Despite the numerous benefits discussed in the literature, TDD is still manual, time-consuming, and labor-intensive, highlighting the need for automated Text-To-Testcase generation.
@@ 2.2 Large Language Models (LLMs)
A language model (LM) is a statistical model trained to predict the next word in a sequence [34], e.g., for the sequence “Paris is the capital of”, it will predict with a very high likelihood that “France” is the next word in the sequence. Large language models are LMs with a substantially large number of weights and parameters and a complex training architecture trained to perform various downstream NLP tasks, e.g., text generation, question answering, and text classification [35]. The recent advancements in LLMs, most notably with chatbot platforms built on LLMs, e.g., OpenAI’s ChatGPT, have led to several applications in software engineering (SE) [36, 37]. These include, among others, requirements engineering automation tasks [38], code generation and completion [39], software vulnerability [40], test generation [41], and software maintenance [36].
@@
Prompting. In generative AI models (built on LLMs), the input is provided as a natural language instruction to elicit a desired response or output [42]. The input is known as a “prompt”. How well a prompt is crafted determines the quality and relevance of the AI’s output. There are different prompting strategies, and one typically needs to experiment with various prompts to elicit accurate responses or outputs from the models [36].
Fine-tuning In the context of Generative AI and ML refers to taking a pre-trained model and further training it on a specific, usually smaller, dataset to adapt it to a particular task or set of tasks. This process is prevalent in applications where a model, initially trained on a diverse and general dataset, needs to be specialized for more specific purposes. This specialized dataset is tailored to meet the particular needs or context of the intended application or task. This approach is beneficial as it leverages the extensive learning from the initial training phase, making the fine-tuning process more efficient than training a new model from scratch on the specialized data. Therefore, fine-tuning enhances the model’s performance and accuracy in tasks requiring a more focused understanding or specialized knowledge [42].
@@ 2.3.1 Automated Test Case Generation
Automated test case generation is an emerging technique that leverages neural networks and other deep learning architectures to automatically create test cases for software applications. This approach aims to enhance the efficiency and effectiveness of software testing by automating the generation of test inputs and expected outputs. In the past few years, researchers proposed automated test case generation approaches such as EvoSuite [8] and Randoop [7] based on different techniques, e.g., random-based approaches, search-based approaches and deep learning-based approaches.
@@ 2.3.2 LLMs for Test Case Generation
Large Language Models (LLMs) have emerged as powerful tools in various fields of natural language understanding and generation, including software engineering [37]. Notably, Hou et al. [37] and Wang et al. [45] found that LLMs have been widely used to support various software testing tasks. For example, Schafer et al. [46] proposed TestPilot, a LLM-based test generation tool for JavaScript that automatically generates unit tests for all API functions in an npm package. Siddiq et al. [47] leveraged Large Language Models for test case generation. Similarly, Yuan et al. [41] and Xie et al. [43] leveraged ChatGPT for test case generation. Moradi Dakhel et al. [48] proposed an approach that utilizes pre-trained Large Language Models in conjunction with mutation testing to generate effective test cases. Yang et al. [49] conducted an empirical study to assess the performance of LLMs in unit test generation. Additionally, Hossain et al. [50] introduced TOGLL, a framework for generating correct and robust test oracles using LLMs.
@@
2.3.3 Effectiveness of Test Case
The effectiveness of test cases in revealing bugs is often measured using mutation testing [51], which evaluates a test suite’s ability to detect small changes to the source code. This aspect is crucial as it reflects the practical utility of the test cases in a real-world testing scenario. The literature consistently indicates that while code coverage is a valuable metric, it does not necessarily correlate strongly with the ability of tests to detect faults. Research has shown that traditional metrics, such as code coverage, do not necessarily correlate strongly with the effectiveness of test cases in identifying software defects. Cai and Lyu [52] suggest that higher coverage does not always equate to higher fault detection rates, especially under varying test profiles. Gopinath et al. [53] examined that while coverage data is considered helpful, it is often insufficient for assessing the suite’s effectiveness in catching critical bugs. Hemmati  [54] provided a critical analysis of different code coverage criteria and their effectiveness in identifying bugs. In light of these findings, our research aims to bridge the gap by not only assessing test cases based on syntax correctness, code coverage, and requirement alignment, but also exploring their capability to detect bugs to understand the effectiveness of test case.